OpenStudy (anonymous):

find the major axis of an ellipse 4x^2+9y^2-8x-18y-23=0 find the major axis of an ellipse 4x^2+9y^2-8x-18y-23=0 @Mathematics

OpenStudy (nikvist):

\[4x^2+9y^2-8x-18y-23=0\]\[4(x^2-2x+1)+9(y^2-2y+1)-36=0\]\[4(x-1)^2+9(y-1)^2=36\]\[\frac{(x-1)^2}{3^2}+\frac{(y-1)^2}{2^2}=1\]\[\mbox{major axis: }y=1\quad,\quad\mbox{length of major axis: }6\]